# Budgeted Spatial Data Acquisition: When Coverage and Connectivity Matter

### v2

@@ 5.2.2 Acceleration Strategies
Index-based Acceleration. We first extend the ball tree [90] to speed up the construction of the spatial dataset graph. Given a set of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\dots,S_{D_{|\mathcal{D}|}}\}$, each dataset can be treated as a ball node. The centroid of the dataset serves as the center of the ball node, and the radius is determined by the maximum distance between the centroid and any cell coordinates within the dataset. This differs from the concept of the center and radius of the dataset graph in Definition 14. We then construct a tree index in a top-down way based on all ball nodes [^1] Please refer to Sec. 6.1 [85] for further details.
@@
BFS-based Acceleration. To speed up finding the center, we adopt a two-round BFS strategy. Specifically, we start a round of BFS by randomly choosing a node $v_{i}$, which enables us to find the node $v_{j}$ that is furthest away from $v_{i}$ (if there are multiple farthest nodes, we randomly choose one). We then take the node $v_{j}$ as the starting node, redo the BFS, and find the node $v_{k}$ farthest from $v_{j}$, where the shortest path distance (defined in Def. 11) from $v_{j}$ to $v_{k}$ is the diameter of the graph, and the center of the path is one of the centers of the graph. Next, we prove the correctness of the above search process in an acyclic graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}},\delta}=(\mathcal{V},\mathcal{E})$.
@@
For the node $v_{j}$ farthest from the given node $v_{i}$, there are three possible cases in total. In the following, we will prove the first two cases are not true, and only the third case holds true.
@@
Case 1: $|\mathcal{SP}(v_{j},v_{k})|>\sigma$. Due to the diameter being the maximum eccentricity of any node in the graph (defined in Def. 13), there cannot be a shortest path such that the distance $|\mathcal{SP}(v_{s},v_{e})|$ is greater than the diameter. Thus, this case is not true.
@@ 6 Experiments
Data Collection Storage (GB) Number of Datasets
BTAA 3.38 3,204
Trackable 0.68 10,000
Identifiable 0.85 10,000
Public 0.54 10,000
OSM 4.07 60,000
@@
6.1 Experimental Setups
Data Collections. We conduct experiments in five real-world spatial data collections. Table 2 shows the statistics of five data collections. The BTAA is sourced from the Big Ten Academic Alliance Geoportal [^2] https://geo.btaa.org/, which offers spatial datasets for the midwestern US, such as bus route, vehicle radar, and census data. The other four collections are public datasets we collected from OpenStreetMap [^3] https://www.openstreetmap.org/traces, which contains real spatial datasets from the world uploaded by users and marked by a tag (e.g., trackable, identifiable, and public). The Trackbale contains the movement tracks of anonymous entities labelled trackable that can be tracked by satellites or sensors, the Identifiable contains the movement tracks of identifiable individuals labelled identifiable, such as vehicles and bikes, and the Public contains the publicly available location datasets labelled public, such as stations and buildings. OSM contains large-scale datasets of three tags simultaneously to verify the scalability of our algorithms.
@@ 6.1 Experimental Setups
    Budget ($B$). We use the ratio of the total prices of all datasets in each data collection as the buyer’s budget.
@@
    Distance threshold ($\delta$) (Def. 6). It specifies the distance requirement for any pair of datasets to be considered to be directly connected.
@@
    Resolution ($\theta$) (Def. 3). It determines the size of unified cells in the grid.
Note that the choice of $\delta$ relies on the specific characteristics of the datasets. E.g., the $\theta$ can be set based on the desired sampling distance in actual scenarios (see Section A.3 in [91]). Specifically, one degree of longitude or latitude is about 111km. If we divide the globe into a $2^{13}\times 2^{13}$ grid, then each cell’s size is about 5km ×2.5km. The $\delta$ can be set by dividing the required distance (e.g., commuting distance) in the actual situation by the cell’s side length. Additionally, our study is orthogonal to the choice of pricing function. Changing pricing functions only impacts the calculation of prices. It will not affect our search algorithm. Here, the price of each spatial dataset is set to the number of covered cells [73, 71, 68].
@@
  • CMC+MC iteratively selects the path with the maximum average coverage (MC) per node on the path despite the marginal gain.
  • CMC+MG iteratively selects the path with the maximum average marginal gain (MG) per node on the path.
@@
Environment. We run all experiments on a 10-core Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHZ processor, with 376G memory. All codes are implemented in Java 1.8, and any method that cannot finish within 24 hours will be terminated. We use Mapv [92] to visualize the returned results in maps. The implementation code and guidelines can be obtained from the GitHub repository [^4] https://github.com/yangwenzhe/BMCC.
6.2 Effectiveness Study
Effectiveness Comparison with Different Budget $B$. Figure 7 compares the performance of the five algorithms with different $B$. We can observe that DSA shows good performance when the budget ratio is 0.001 and 0.005. However, as the budget gradually increases, the performance of the DSA becomes worse than the other four algorithms. This is because the DSA is highly dependent on the selection of initial nodes. Additionally, since the graph we input is composed of multiple subgraphs, the DSA may select the initial node belonging to one of the subgraphs and get stuck in a locally optimal solution.
Moreover, we observe that as $B$ increases, the solutions obtained by the DPSA cover at most 68% times coverage than CMC+MC and 21% times coverage than CMC+MG. This indicates the superior effectiveness of DPSA over the previously proposed algorithm. Additionally, we can find that no matter how the budget changes, the solutions obtained by DPSA and DPSA+BA are within 1% of the error. This demonstrates the effectiveness of our proposed BFS-based acceleration strategy for quickly finding the center. Notably, the experimental results of DPSA on OSM are not presented due to the algorithm’s inability to complete within 24 hours. This is attributed to the time-consuming process of finding the center in DPSA.
Effectiveness Comparison with Different Threshold $\delta$. Figure 8 compares the performance of the five algorithms with different $\delta$. Firstly, as $\delta$ increases, the spatial coverage found by the five algorithms under the same budget gradually increases. This is because a higher $\delta$ means that more pairs of datasets satisfy the directly connected relation, resulting in a higher average degree in the graph, as shown in Table 4. Thus, the spatial coverage in the solution found by the algorithm also increases. Additionally, we can observe that the DPSA consistently demonstrates robust performance irrespective of $\delta$ variations, achieving at most 63% and 35% times spatial coverage against the CMC+MG and the CMC+MC. This indicates that our algorithm is effective compared to the two variants.
Effectiveness Comparison with Different Resolution $\theta$. Figure 9 compares the performance of the five algorithms with different $\theta$. We can see that as the $\theta$ increases, the spatial coverage is gradually increasing. This can be explained by the fact that in the stage of generating the cell-based dataset, we divide the entire space into a $2^{\theta}\times 2^{\theta}$ grid and generate the cell-based dataset for each raw dataset. Thus, the spatial coverage of each cell-based dataset increases as $\theta$ increases, leading to larger spatial coverage found by the five algorithms. Additionally, the DPSA and DPSA+BA can achieve at most 58% and 38% times spatial coverage against the CMC+MG and CMC+MC, respectively, which also verifies the effectiveness of our proposed algorithms.
@@ 6.2 Effectiveness Study
6.3 Efficiency Study
Efficiency Comparison with Different Budget $B$. Figure 10 compares the running time of the five algorithms with different $B$. We have five observations:
(1) The DSA always has the shortest running time compared to the other algorithms, completing the search in a few milliseconds. Combined with the results shown in Figure 7, we can see that the DSA offers the best trade-off in terms of speed and accuracy when the budget ratio does not exceed 0.005.
(3) Although DPSA can find a better solution when the budget ratio exceeds 0.05, it exhibits significantly slower performance, typically by one or two orders of magnitude, compared to CMC+MC and CMC+MG. This is because the process of finding the center node takes significant time in DPSA. Additionally, the DPSA+BA can achieve up to at most 89% times speedups over CMC+MG, providing compelling evidence to support the proposition that utilizing the BFS-based acceleration strategy for expedited center node identification is highly efficient.
(5) When $B$ is small, CMC+MG runs slightly faster than DPSA. However, as $B$ increases, CMC+MG takes longer execution times compared to DPSA. This is because when $B$ is small, the number of affordable datasets is small (i.e., the algorithm has fewer iterations), making the running time of finding the center more important in the overall algorithm. CMC+MG does not need to find a center, so the algorithm is slightly faster with a smaller $B$. However, as $B$ increases, the number of iterations also increases, and CMC+MG needs to traverse more paths than DPSA, resulting in a longer running time for CMC+MG. Furthermore, combined with Figure 7, we can see that the result set found by the CMC+MG is worse than DPSA+BA, which indicates that the two maximization strategies adopted in our algorithm are effective and robust.
Efficiency Comparison with Different Threshold $\delta$. Figure 11 compares the running time with different $\delta$. We can observe that as $\delta$ increases, the running time of the five algorithms gradually increases. Among them, DSA has the shortest running time and the smallest increase, while DPSA has the maximum increase. This is because the time taken to find the center in the DPSA increases significantly as $\delta$ increases. Another interesting phenomenon is that apart from DPSA and DSA, CMC+MG has the slowest running time as $\delta$ increases, followed by DPSA+BA, and CMC+MC is the fastest. This is because CMC+MG builds the path from the root node to all other nodes at each iteration, while DPSA+BA only considers the path from the root node to the leaf node. Combined with Figure 7, we can see that although the running time of CMC+MC is shorter than that of DPSA+BA and CMC+MG, the solution found is less effective because it does not consider the marginal gain.
Efficiency Comparison with Different Resolution $\theta$. Figure 12 compares the running time of the five algorithms with different $\theta$. We have an interesting observation from this figure. As the $\theta$ increases, only the running time of DPSA decreases, while the running time of the other four algorithms increases gradually. This is because as the $\theta$ increases, the entire space is divided into finer granularity, resulting in a higher spatial coverage in each cell-based dataset. Furthermore, the finer granularity of each dataset contributes to an increase in spatial dataset distance between pairs of datasets, causing more node pairs that cannot be directly connected. This, in turn, leads to an increase in the number of subgraphs and a decrease in the average degree of subgraphs, as shown in Table 4. Thus, the time taken to find the center in DPSA decreases significantly, resulting in a decrease in the overall running time of DPSA.
@@ 6.4 Visual Analysis
For visualization purposes, we show 2000 individual trajectories located at Washington from BTAA, as shown in Figure 13(a). The solutions found by five algorithms with the default settings are shown in Figures 13(b)–(f). As shown in Figure 13(b), We can clearly see that the datasets obtained by the DSA exhibit lower coverage, whereas the datasets obtained by the CMC+MG demonstrate higher coverage compared to CMC+MC and DSA when the budget is 0,1. This reveals that the performance of the DSA becomes worse than other algorithms when the budget is larger due to the decrease in the approximation ratio. Additionally, the datasets found by CMC+MC are only distributed on one side of the entire space, which has a one-sidedness. In contrast, we observe that the datasets obtained by the DPSA and DPSA+BA are distributed over the entire space. This shows that our algorithm has better spatial coverage compared to baselines, which validates the effectiveness of our proposed algorithms. For data buyers who do not have the budget to acquire all datasets, this solution is a good representation of the dataset distribution in the entire data marketplace.

### v3

@@ 5.2.2 Acceleration Strategies
Index-based Acceleration. We first extend the ball tree [79] to speed up the construction of the spatial dataset graph. Given a set of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\dots,S_{D_{|\mathcal{D}|}}\}$, each dataset can be treated as a ball node. The centroid of the dataset (i.e., the mean of all cell coordinates within the dataset) serves as the center of the ball node, and the radius is determined by the maximum distance between the centroid and any cell coordinates within the dataset. This differs from the concept of the center and radius of the dataset graph in Definition 11. We then construct a tree index in a top-down way based on all ball nodes [74].
@@
BFS-based Acceleration. To speed up finding the center, we adopt a two-round BFS strategy. Specifically, we start a round of BFS by randomly choosing a node $v_{i}$, which enables us to find the node $v_{j}$ that is furthest away from $v_{i}$ (if there are multiple farthest nodes, we randomly choose one). We then take the node $v_{j}$ as the starting node, redo the BFS, and find the node $v_{k}$ farthest from $v_{j}$, where the shortest path distance (defined in Def. 8) from $v_{j}$ to $v_{k}$ is the diameter of the graph, and the center of the path is one of the centers of the graph. Next, we prove the correctness of the above search process in an acyclic graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}},\delta}=(\mathcal{V},\mathcal{E})$.
@@
For the node $v_{j}$ farthest from the given node $v_{i}$, there are three possible cases in total. In the following, we will prove that the first two cases are not true, and only the third case holds true.
@@
Case 1: $|\mathcal{SP}(v_{j},v_{k})|>\sigma$. Due to the diameter being the maximum eccentricity of any node in the graph (defined in Def. 10), there cannot be a shortest path such that the distance $|\mathcal{SP}(v_{s},v_{e})|$ is greater than the diameter. Thus, this case is not true.
@@ 6 Experiments
Data Collection Storage (GB) Number of Datasets Number of Points Coordinates Range
BTAA 3.38 3,204 96,788,280 [$(-179^{\circ}.46^{\prime},-87^{\circ}.70^{\prime}),(179^{\circ}.98^{\prime},71^{\circ}.40^{\prime})$]
Trackable 0.68 10,000 40,442,724 [$(-53^{\circ}.16^{\prime},-89^{\circ}.99^{\prime}),(69^{\circ}.70^{\prime},88^{\circ}.21^{\prime})$]
Identifiable 0.85 10,000 67,688,576 [$(-89^{\circ}.68^{\prime},-89^{\circ}.99^{\prime}),(88^{\circ}.88^{\prime},89^{\circ}.99^{\prime})$]
Public 0.54 10,000 32,407,907 [$(-54^{\circ}.85^{\prime},-89^{\circ}.99^{\prime}),(79^{\circ}.44^{\prime},76^{\circ}.38^{\prime})$]
OSM 4.07 60,000 207,879,164 [$(-68^{\circ}.25^{\prime},-89^{\circ}.99^{\prime}),(70^{\circ}.06^{\prime},89^{\circ}.99^{\prime})$]
6.1 Experimental Setups
Data Collections. We conduct experiments in five real-world spatial data collections. Table 2 shows the statistics of five data collections, which covers the storage, the number of datasets, the number of points, and the range of coordinates. The BTAA is sourced from the Big Ten Academic Alliance Geoportal, which offers spatial datasets for the midwestern US, such as bus routes, vehicle radar, and census data. The other four collections are public datasets we collected from OpenStreetMap, which contain real spatial datasets from the world uploaded by users and marked by a tag (e.g., trackable, identifiable, and public). The Trackbale contains the movement tracks of anonymous entities labelled trackable that can be tracked by satellites or sensors, the Identifiable contains the movement tracks of identifiable individuals labelled identifiable, such as vehicles and bikes, and the Public contains the publicly available location datasets labelled public, such as stations and buildings. OSM contains large-scale datasets of three tags simultaneously to verify the scalability of our algorithms. Fig. 7 shows the distribution of dataset scale in five dataset collections. We can observe that the dataset distribution is relatively uniform across the different scales, ensuring a comprehensive evaluation across various scenarios.
@@ 6.1 Experimental Setups
$m$: number of datasets {0.2, 0.4, 0.6, 0.8, 1}×$\times|\mathcal{D}|$
Figure 7: Dataset distributions in five data collections.
@@
    Budget ($B$). It defines the total allowable budget for selecting datasets.
@@
    Distance threshold ($\delta$). It specifies the distance requirement for any pair of datasets to be considered directly connected.
@@
    Resolution ($\theta$). It determines the size of uniform cells in the grid. The larger the value of $\theta$, the smaller the cell size, and vice versa.
  • $\bullet$
    Scale ($m$). It specifies the scale of the dataset in each dataset collection used to verify the scalability of our algorithms.
Firstly, we use the ratio of the total prices of all datasets in each data collection as the buyer’s budget. Moreover, the parameter $\theta$ can be set based on the desired sampling distance in actual scenarios. Specifically, one degree of longitude or latitude is about 111km. If we divide the globe into a $2^{13}\times 2^{13}$ grid, then each cell’s size is about 5km ×2.5km. Furthermore, we vary the dataset scale from 0.2×$\times|\mathcal{D}|$ to 1.0×$\times|\mathcal{D}|$ to verify the scalability of our algorithms. Finally, the $\delta$ can be set by dividing the required distance (e.g., commuting distance) in the actual situation by the cell’s side length. It is worth noting that our study is orthogonal to the choice of pricing function. Changing pricing functions only impacts the calculation of prices and does not impact our search algorithm. Here, the price of each spatial dataset is set to the number of covered cells [40, 38, 33].
@@
  • CMC+MC selects an arbitrary node to build a BFS tree and iteratively selects the path with the maximum average coverage (MC) per node on the path despite the marginal gain.
  • CMC+MG selects an arbitrary node to build a BFS tree and iteratively selects the path with the maximum average marginal gain (MG) per node on the path.
@@
Environment. We run all experiments on a 10-core Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHZ processor, with 376G memory. All codes are implemented in Java 1.8, and any method that cannot finish within 24 hours will be terminated.
6.2 Effectiveness Study
Effectiveness Comparison with Different Budget $B$. Fig. 8 compares the performance of the five algorithms with different $B$. We can observe that DSA shows good performance when the budget ratio is 0.001 and 0.005. However, as the budget gradually increases, the performance of DSA becomes worse than the other four algorithms. This is because DSA is highly dependent on the selection of initial nodes. Additionally, since the graph we input is composed of multiple subgraphs, DSA may select the initial node belonging to one of the subgraphs and get stuck in a locally optimal solution.
Moreover, we observe that with increasing $B$, the solutions obtained by DPSA cover at most 68% times coverage than CMC+MC and 21% times coverage than CMC+MG. This indicates the superior effectiveness of DPSA over the previously proposed algorithm. Additionally, we can find that no matter how the budget changes, the solutions obtained by DPSA and DPSA+BA are within 1% of the error. This demonstrates the effectiveness of our proposed BFS-based acceleration strategy for quickly finding the center. Notably, the experimental results of DPSA on OSM are not presented due to the algorithm’s inability to complete within 24 hours. This is attributed to the time-consuming process of finding the center in DPSA.
Effectiveness Comparison with Different Threshold $\delta$. Fig. 9 compares the performance of the five algorithms with different $\delta$. Firstly, as $\delta$ increases, the spatial coverage found by the five algorithms under the same budget gradually increases. This is because a higher $\delta$ means that more pairs of datasets satisfy the directly connected relation, resulting in a higher average degree in the graph, as shown in Table 4. Thus, the spatial coverage in the solution found by the algorithm also increases. Additionally, we can observe that DPSA consistently demonstrates robust performance irrespective of $\delta$ variations, achieving at most 63% and 35% times spatial coverage against CMC+MG and CMC+MC. This indicates that our algorithm is effective compared to the two variants.
Effectiveness Comparison with Different Resolution $\theta$. Fig. 10 compares the performance of the five algorithms with different $\theta$. We can see that as the $\theta$ increases, the spatial coverage is gradually increasing. This can be explained by the fact that in the stage of generating the cell-based dataset, we divide the entire space into a $2^{\theta}\times 2^{\theta}$ grid and generate the cell-based dataset for each raw dataset. Thus, the spatial coverage of each cell-based dataset increases as $\theta$ increases, leading to larger spatial coverage found by the five algorithms. Additionally, DPSA and DPSA+BA can achieve at most 58% and 38% times spatial coverage against CMC+MG and CMC+MC, respectively, which also verifies the effectiveness of our proposed algorithms.
6.3 Efficiency Study
Efficiency Comparison with Different Budget $B$. Fig. 11 compares the running time of the five algorithms with different $B$, with execution time measured in milliseconds (ms). We have five observations:
(1) DSA always has the shortest running time compared to the other four algorithms, completing the search in a few milliseconds. Combined with the experimental results shown in Figs. 8 and 11, we can see that DSA offers the best trade-off in terms of speed and accuracy when the budget ratio does not exceed 0.005.
(3) Although DPSA can find a better solution when the budget ratio exceeds 0.05, it exhibits significantly slower performance, typically by one or two orders of magnitude, compared to CMC+MC and CMC+MG. This is because the process of finding the center node takes significant time in DPSA. Additionally, DPSA+BA can achieve up to at most 89% times speedups over CMC+MG, providing compelling evidence to support the proposition that utilizing the BFS-based acceleration strategy for expedited center node identification is highly efficient.
(5) When $B$ is small, CMC+MG runs slightly faster than DPSA. However, as $B$ increases, CMC+MG takes longer execution times compared to DPSA. This is because when $B$ is small, the number of affordable datasets is small (i.e., the algorithm has fewer iterations), making the running time of finding the center more important in the overall algorithm. CMC+MG does not need to find a center, so the algorithm is slightly faster with a smaller $B$. However, as $B$ increases, the number of iterations also increases, and CMC+MG needs to traverse more paths than DPSA, resulting in a longer running time for CMC+MG. Furthermore, combined with Fig. 8, we can see that the result set found by CMC+MG is worse than DPSA+BA, which indicates that the two maximization strategies adopted in our algorithm are effective and robust.
Efficiency Comparison with Different Threshold $\delta$. As shown in Fig. 12, we can observe that as $\delta$ increases, the running time of the five algorithms gradually increases. Among them, DSA has the shortest running time and the smallest increase, while DPSA has the maximum increase. This is because the time taken to find the center in DPSA increases significantly as $\delta$ increases. Another interesting phenomenon is that apart from DPSA and DSA, CMC+MG has the slowest running time as $\delta$ increases, followed by DPSA+BA, and CMC+MC is the fastest. This is because CMC+MG builds the path from the root node to all other nodes at each iteration, while DPSA+BA only considers the path from the root node to the leaf node. Combined with Fig. 8, we can see that although the running time of CMC+MC is shorter than that of DPSA+BA and CMC+MG, the solution found is less effective because it does not consider the marginal gain.
Efficiency Comparison with Different Resolution $\theta$. Fig. 13 compares the running time of the five algorithms with different $\theta$. We have an interesting observation from this figure. As the $\theta$ increases, only the running time of DPSA decreases, while the running time of the other four algorithms increases gradually. This is because as the $\theta$ increases, the entire space is divided into finer granularity, resulting in a higher spatial coverage in each cell-based dataset. Furthermore, the finer granularity of each dataset contributes to an increase in spatial dataset distance between pairs of datasets, causing more node pairs that cannot be directly connected. This, in turn, leads to an increase in the number of subgraphs and a decrease in the average degree of subgraphs, as shown in Table 4. Thus, the time taken to find the center in DPSA decreases significantly, resulting in a decrease in the overall running time of DPSA.
6.4 Scalability
The objective of this experiment is to evaluate the scalability of our algorithms and two comparison algorithms when we increase the dataset size. Fig. 14 shows the results of five algorithms’ spatial coverage and running time, respectively, when we vary the dataset size from 0.2×$\times|\mathcal{D}|$ to 1.0×$\times|\mathcal{D}|$ and set the other parameters as defaults. From the left of Fig. 14, we can observe that the spatial coverage of five algorithms slightly increases as the dataset scale increases. This is because, as the dataset scale increases, each dataset has more potential connectable datasets, leading to a higher degree of nodes in the dataset graph. Therefore, under the given budget, the algorithm is able to select more datasets that contribute to improving spatial coverage, thus enhancing the overall coverage effect. However, no matter how the size of the dataset increases, our DPSA and DPSA+BA can still maintain a relatively larger spatial coverage than other comparison algorithms, which exhibit good scalability characteristics.
From the right of Fig. 14, we can see that the running time of five algorithms slightly increases as the dataset scale increases. This is because the algorithms need more time to process the increased number of connectable nodes. Among the five algorithms, DSA consistently exhibits the shortest running time, completing the search in just a few milliseconds, which demonstrates its strong scalability. However, due to the relatively large budget set by default in this scalability experiment, the approximation ratio of DSA decreases, resulting in its spatial coverage performance being worse than that of the other comparison algorithms. This observation is consistent with our theoretical analysis. In addition to DSA, we can observe that DPSA+BA has a faster search speed than other methods, which indicates that DPSA+BA can efficiently handle larger datasets and maintain high performance even as the size of the data increases.
Figure 14: Comparison of five algorithms’ spatial coverage and running time with dataset scale $m$ increasing.
@@ 6.5 Visual Analysis
For visualization purposes, we show 2000 individual trajectories located at Washington from BTAA, as shown in Fig. 15(a). The solutions found by five algorithms with the default settings are shown in Figs. 15(b)–(f). As shown in Fig. 15(b), We can clearly see that the datasets obtained by DSA exhibit lower coverage, whereas the datasets obtained by CMC+MG demonstrate higher coverage compared to CMC+MC and DSA when the budget is 0,1. This reveals that the performance of DSA becomes worse than other algorithms when the budget is larger due to the decrease in the approximation ratio. Additionally, the datasets found by CMC+MC are only distributed on one side of the entire space, which has a one-sidedness. In contrast, we observe that the datasets obtained by DPSA and DPSA+BA are distributed over the entire space. This shows that our algorithm has better spatial coverage compared to baselines, which validates the effectiveness of our proposed algorithms. For data buyers who do not have the budget to acquire all datasets, this solution is a good representation of the dataset distribution in the entire data marketplace.